OpenStudy (anonymous):

When it hatches from its egg, the shell of a certain crab is 1cm across.? When fully grown the shell is approximately 10cm across. Each new shell is one-third bigger than the previous one. How many shells does a fully grown crab have during its life

OpenStudy (anonymous):

Hatching dimension of the shell,a0 = 1cm The first shall dimension , a1=1+1*(1/3) = 4/3 cm The 2nd shell dimension a2=a1*+a1*(1/3) =(4/3)^2 cm The 3rd shell dimension ,a3= a2+a2(1/3) = a2(4/3) =(4/3)^3 cm Therefore, the nth shell dimension , an =(4/3)^n cm If nth shell is full life size , then (4/3)^n = 10. Taking logarithms , we can solve for n. n log(4/3) = log10 n= log10/log(4/3) = 8.00392278 . Therefore, an , is the 8th shell whose dimension is nearly 10cm. By actual calculation also we can see the dimensions of the shells at varoius stages. a1=1.3333 cm a2=1.7777 cm a3=2.3704 cm a4=3.1605 cm a5=4.2140 cm a6=5.6187 cm a7=7.4915 cm a8=9.9887 cmis the 8th shell. a9=13.3183 >10 . So not possible So, a8=9.887cm is the dimension. a9==13.3183cm is not possible as it is more than 10cm.